OpenStudy (anonymous):

How to solve for x: y=(x-2) (x+2) (x^2+2) Derived from: -x^4 + 2x^2 + y +8 = 0

OpenStudy (anonymous):

Well if y=0, then one of the factors must equal zero too. Taking in all possibilities, we have \[x-2=0 \text{ so } x=2\]\[x+2=0 \text{ so } x=-2\]\[x^2+2=0 \text{ has no real solutions}\]Therefore\[x=\left\{-2,2\right\}\]For more solutions to textbook problems, be sure to check out http://www.slader.com/s/eWFrZXlnbGVl . I submit a lot of content there as well.

OpenStudy (anonymous):

you have given only 1 equation but there is 2 unknown quantity are there x and y so x can be find in terms of y only and i want to ask from yakeyglee why we will take y as 0